OpenStudy (anonymous):

graph l x-3l = y what steps do i do to find out what points to plot

OpenStudy (amistre64):

x=|y= ------ | | | |

OpenStudy (amistre64):

make a table of points...

OpenStudy (anonymous):

|x-3|=y means that two equations are there 1) x-3=y 2)x-3=-y; then solve as you find points in equation 1 & 2 put some points of x and evaluate corresponding y

OpenStudy (amistre64):

when x=0; y=3 when x=-1; y=4 when 3=1; y=2...etc

OpenStudy (amistre64):

x=1...not 3=1 lol

OpenStudy (anonymous):

say x=0 gives you -3 for equ 1 and 3 for equation 2. and so on

OpenStudy (amistre64):

splitting it up complicates matters; makes you want to form an x

OpenStudy (anonymous):

haha a typo isnt as bad as feeling stupid for not being able to figure out something such as this. I am just fried. Learning a month of algebra in one week

OpenStudy (amistre64):

when x=5; y=2 and y=-2? nah.... when x=5; y=2

OpenStudy (amistre64):

in fact y is never gonna be a negative number

OpenStudy (anonymous):

i give a plot of the graph...

OpenStudy (anonymous):

Thanks For the help guys. i appreciate it!

OpenStudy (amistre64):

that aint it lol.... its y=|x| shifted to the right by 3

OpenStudy (amistre64):

that one you drew up is y=|x|-3

OpenStudy (amistre64):

just plug in numbers for x and tabulate the y from there; plain and simple